### v3

# Adversarial control of synchronization in complex oscillator networks

Abstract
This study investigates perturbation strategies inspired by adversarial attack principles from deep learning, designed to control synchronization dynamics through strategically crafted weak perturbations. We propose a gradient-based optimization method that identifies small phase perturbations to dramatically enhance or suppress collective synchronization in Kuramoto oscillator networks. Our approach formulates synchronization control as an optimization problem, computing gradients of the order parameter with respect to oscillator phases to determine optimal perturbation directions. Results demonstrate that extremely small phase perturbations applied to network oscillators can achieve significant synchronization control across diverse network architectures. Our analysis reveals that synchronization enhancement is achievable across various network sizes, while synchronization suppression becomes particularly effective in larger networks, with effectiveness scaling favorably with network size. The method is systematically validated on canonical model networks including scale-free and small-world topologies, and real-world networks representing power grids and brain connectivity patterns. This adversarial framework represents a novel paradigm for synchronization management by introducing deep learning concepts to networked dynamical systems.
Synchronization emerges naturally in many networked systems, where individual components coordinate their behavior to act collectively. Controlling such synchronization in complex networks has been challenging, typically requiring substantial interventions. This study introduces a novel approach inspired by adversarial attack principles from artificial intelligence, where tiny, strategically designed perturbations can dramatically alter system behavior. We applied this concept to networks of oscillating systems and discovered that extremely small adjustments can either enhance or suppress synchronization across the entire network. These weak interventions become particularly effective in larger networks, suggesting excellent scalability for real-world applications. We validated this approach on various network types including power grids and brain networks, demonstrating potential applications for infrastructure stability and managing pathological brain activity. This research offers a new paradigm for controlling complex networked systems with weak intervention.
I Introduction
Synchronization phenomena in complex networks of coupled oscillators are ubiquitous in nature, spanning from the coordinated flashing of fireflies and cardiac pacemaker cells to the stability of power grids and brain network dynamics Arenas et al. (2008); Suykens and Osipov (2008); Dörfler and Bullo (2014). The Kuramoto model has emerged as a fundamental paradigm for understanding these collective behaviors by focusing on the phase dynamics of weakly coupled oscillators Acebrón et al. (2005); Rodrigues et al. (2016). Despite its apparent simplicity, the model captures essential features of synchronization transitions and has become an invaluable theoretical framework for investigating synchronization control strategies across diverse scientific disciplines.
Traditional synchronization control approaches in Kuramoto oscillator networks primarily target the system’s intrinsic parameters. Network topology significantly influences synchronization dynamics, with scale-free networks promoting synchronization through highly connected hub nodes Moreno and Pacheco (2004); Peron et al. (2019); Medvedev and Tang (2020) and small-world networks enhancing synchronization via shortcuts that reduce effective path lengths between distant oscillators Hong, Choi, and Kim (2002); Grabow et al. (2010); Nishikawa et al. (2003). Based on this understanding, topology modification approaches strategically add or remove connections or implement adaptive rewiring to alter critical coupling thresholds Zhou and Kurths (2006); Chen et al. (2008); Zhang and Xiao (2014); Pinto and Saa (2015); Papadopoulos et al. (2017); Ghorban, Hesaam, and Sarbazi-Azad (2022); Ruangkriengsin and Porter (2023); Song et al. (2024), though such structural modifications require physical access to network connections in biological or infrastructure systems. Natural frequency manipulation exploits the principle that narrow frequency distributions promote synchronization Acebrón and Spigler (1998); Taylor, Ott, and Restrepo (2010); Gil (2021), while maintaining non-natural frequencies requires external energy sources. Furthermore, coupling strength adjustment represents another approach Ren and Zhao (2007); Ha, Noh, and Park (2016); Cho, Dong, and Ha (2023), with specific connection weight distributions such as frequency-dependent coupling influencing synchronization Wang and Li (2011), though this can lead to complex dynamical behaviors Rossi et al. (2023). Optimization approaches for natural frequencies or link weights have also been proposed Menara et al. (2022), with computational requirements scaling with network size.
External intervention strategies control synchronization without permanently altering system characteristics. Pinning control applies targeted signals to strategically selected nodes to drive network-wide synchronization Wang and Chen (2002); Sorrentino et al. (2007); Xiang et al. (2007); Yang et al. (2010); Lin et al. (2016), with performance depending on appropriate node selection in heterogeneous networks. Time-delayed feedback and global feedback mechanisms (such as mean-field feedback) can promote or suppress synchronization by monitoring the collective state of the system and adjusting control signals accordingly Yeung and Strogatz (1999); Rosenblum and Pikovsky (2004); Taher, Olmi, and Schöll (2019); Franci et al. (2012); Ozawa and Kori (2021), requiring optimization of delay times and feedback strengths, with performance being sensitive to these parameter settings. Noise-induced synchronization exploits the principle that appropriately tuned random perturbations can enhance collective behavior through mechanisms such as stochastic resonance and stochastic synchronization Kawamura, Nakao, and Kuramoto (2007); Esfahani, Shahbazi, and Samani (2012); Lai and Porter (2013); Tyloo, Delabays, and Jacquod (2019), though the stochastic nature of this approach introduces inherent limitations for deterministic control applications. The sensitivity of networked oscillators to strategic perturbations has been further demonstrated in studies of network vulnerabilities, where targeted interventions at individual nodes can produce significant changes in global synchronization behavior Tyloo (2023).
While these approaches have made valuable contributions, the above challenges remain. To address these limitations, we explore an alternative strategy inspired by adversarial attack principles from deep learning. Adversarial attacks involve strategically crafted small perturbations added to inputs that can dramatically alter the outputs of deep neural network models, exploiting inherent system vulnerabilities Yuan et al. (2019); Hirano and Takemoto (2020); Hirano, Koga, and Takemoto (2020); Hirano, Minagi, and Takemoto (2021); Chakraborty et al. (2021); Akhtar et al. (2021); Koga and Takemoto (2022). This concept has recently been extended to dynamics on networks in other domains, specifically applied to voter model dynamics and prisoner’s dilemma games on networks, where small strategic perturbations can significantly alter collective behaviors Chiyomaru and Takemoto (2022, 2023); Mizutaka (2023); Takemoto (2024); Ninomiya et al. (2025). We introduce this adversarial framework to synchronization control in Kuramoto oscillator networks, where the goal is to identify small perturbations to the networked dynamics that can disrupt or enhance synchronization by exploiting the system’s intrinsic sensitivities. This approach offers unique characteristics: requiring only weak perturbations to phases while leveraging optimization principles to achieve large-scale effects on collective behavior. By treating synchronization control as an adversarial problem, we introduce a novel theoretical framework that complements existing approaches and provides new insights into the fundamental mechanisms of synchronization manipulation. While our method draws inspiration from adversarial attacks in deep learning, specifically the principle of using small perturbations to achieve large effects, it is fundamentally a gradient-based control strategy that requires real-time access to system states rather than a machine learning technique. We emphasize that our approach represents a conceptual transfer of the adversarial principle rather than establishing direct mathematical analogies between neural networks and oscillator networks.
In this work, we develop and analyze adversarial attack strategies specifically designed for Kuramoto oscillator networks to achieve synchronization control through phase perturbations. We formulate the adversarial attack as an optimization problem that identifies near-minimal perturbations to oscillator phases that can maximally disrupt or enhance network synchronization. Our approach combines gradient-based optimization techniques with network analysis to systematically exploit the vulnerabilities inherent in the oscillator networks. We demonstrate that synchronization can indeed be effectively controlled through this adversarial framework via extensive numerical simulations on both model networks and real-world networks. Moreover, we investigate and discuss the effects of network size, average node degree, and network connectivity patterns on the outcomes of adversarial attacks.
II Model
We consider the Kuramoto model on networks Moreno and Pacheco (2004); Arenas et al. (2008); Dörfler and Bullo (2014); Rodrigues et al. (2016); Peron et al. (2019), where the time evolution of the phases () of coupled oscillators on a given network is described by
where represents the natural frequency of oscillator , and denotes the coupling strength that characterizes the strength of interactions between oscillators. For simplicity, we consider unweighted and undirected networks. Thus, if there exists a link between oscillators (nodes) and , and otherwise.
Alternative formulations using degree normalized coupling strength Arenas et al. (2008); Rodrigues et al. (2016), where is the degree of node , do not affect our main conclusions, as will be shown when we develop our gradient based attack methodology.
To quantify the degree of synchronization in the network, we define the complex order parameter
where represents the synchronization strength and denotes the average phase. The synchronization strength ranges from 0 to 1, with indicating strong synchronization (all oscillators have similar phases) and indicating an incoherent state (oscillator phases are uniformly distributed).
III Adversarial attcks
Here, we apply the concept of adversarial attacks to promote or suppress synchronization in Kuramoto oscillator networks. Specifically, we use the order parameter as the objective function and apply a gradient descent method to adjust each oscillator’s phase along the gradient direction, thereby guiding the entire system toward a synchronized state.
The gradient-based approach is formulated as
where the gradient is derived as
Here, represents the average phase calculated just before applying the perturbation. Physically, this gradient indicates the direction that brings each oscillator’s phase closer to the average phase, thereby promoting synchronization.
However, obtaining exact phase values and controlling perturbation strength can be challenging. To avoid these limitations, following the principle of the fast gradient sign method Goodfellow, Shlens, and Szegedy (2015), we consider using only the sign of the gradient. Specifically, we add perturbations to each oscillator’s phase at regular time intervals as follows:
When , the adjustment occurs in the opposite direction of the gradient, actively suppressing synchronization.
The parameter controls the perturbation strength, where determines the magnitude of phase perturbations applied to the network.
To compare the performance of the adversarial perturbations with random controls, we also consider random perturbations with the same perturbation strength applied at the same time intervals : , where is a random variable uniformly sampled from the set .
IV Simulations
IV.1 Network models and simulation setup
Our numerical analysis examines the effectiveness of adversarial attacks on synchronization dynamics using three well-established network topologies that capture different structural properties found in real-world systems.
We employ networks consisting of oscillators with an average degree , unless stated otherwise. Robustness tests confirm that varying network size () or average degree ( does not alter the qualitative behavior. Each oscillator’s intrinsic frequency follows a standard normal distribution , while initial phases are randomly assigned from a uniform distribution over .
Three network architectures are investigated: Erdős–Rényi (ER) random networks Albert and Barabási (2002); Takemoto and Oosawa (2012) serve as our baseline topology, where edges are placed between randomly chosen node pairs, yielding Poisson-distributed degrees with mean . To capture heterogeneous connectivity patterns, we also examine Barabási–Albert (BA) scale-free networks Albert and Barabási (2002); Barabási and Albert (1999) constructed through preferential attachment, where new nodes connect to existing nodes, producing power-law degree distributions with . Additionally, Watts–Strogatz (WS) small-world networks Watts and Strogatz (1998) are analyzed, which begin as regular lattices where each node connects to nearest neighbors, then edges are rewired with probability of 0.05 Hirano and Takemoto (2019); Chiyomaru and Takemoto (2022); Takemoto (2024), creating networks with enhanced clustering compared to random graphs.
The dynamics are integrated until the system reaches equilibrium, and the order parameter is measured in this steady state. The time evolution of is computed for each realization, and statistical reliability is ensured by averaging results across 100 independent trials, each with different network topologies and initial phase configurations.
IV.2 Adversarial perturbation effects on synchronization transitions
Figure 1 shows the relationship between the order parameter and coupling strength for different perturbation parameters across three network topologies. The most striking observation is that even small perturbation strengths can significantly alter synchronization transitions regardless of network topology. Positive values (red curves) consistently promote synchronization by shifting the transition to lower coupling strengths, while negative values (blue curves) suppress synchronization, requiring higher coupling strengths to achieve the same level of synchronization across all network types. This demonstrates the remarkable effectiveness of adversarial perturbations in controlling synchronization dynamics with weak intervention strength.
While the fundamental ability to manipulate synchronization transitions is preserved across different network architectures, some topology-dependent variations are observed. BA networks (Fig. 1b) show somewhat reduced susceptibility to synchronization suppression compared to ER networks (Fig. 1a), likely due to their scale-free structure with highly connected hubs that facilitate synchronization. Conversely, WS networks (Fig. 1c) exhibit particularly pronounced responses to positive perturbations. Although WS networks with their lattice-like local structure are inherently difficult to synchronize, adversarial interventions can dramatically enhance their synchronization capability.
IV.3 Perturbation strength effects
To investigate the relationship between perturbation strength and synchronization control more systematically, Figure 2 examines the order parameter as a function of perturbation strength for fixed coupling strength across different network topologies and intervention intervals . The coupling strength is chosen to ensure appropriate baseline conditions: for synchronization enhancement (upper panels), we select values that yield in the unperturbed state, while for synchronization suppression (lower panels), we choose values corresponding to . However, due to the inherent difficulty of synchronization in WS networks, we empirically set for this topology to achieve meaningful perturbation effects.
The results demonstrate a striking sensitivity of synchronization dynamics to perturbation strength. For synchronization enhancement (upper panels), increases sharply with even small positive values, showing that weak perturbations can yield substantial synchronization promotion. Similarly, for synchronization suppression (lower panels), decreases rapidly as the magnitude of negative increases, indicating that small perturbation strengths are sufficient to significantly disrupt collective dynamics. This sharp dependence on underscores the remarkable efficiency of adversarial attacks in manipulating synchronization with weak intervention.
The intervention interval influences the steepness of these transitions. Shorter intervals () produce the sharpest – curves, while longer intervals () result in more gradual changes. This reflects the expected reduction in control effectiveness with less frequent interventions.
The comparison with random perturbations (crosses) reveals the superiority of adversarial attacks. Random perturbations show minimal variation in regardless of perturbation strength, remaining close to the unperturbed baseline. This stark contrast highlights that the effectiveness of adversarial perturbations stems from their strategic targeting rather than mere noise addition.
Network-dependent variations are also evident. WS networks exhibit the strongest responses to both enhancement and suppression attempts, while BA networks show the most resistance to perturbation effects, particularly for synchronization suppression.
IV.4 Network size effects
To examine how network size influences the effectiveness of adversarial attacks, Figure 3 shows the order parameter as a function of network size using the same coupling strength settings as Fig. 2 but with the intervention interval fixed at . The results reveal distinctly different behaviors between synchronization enhancement and suppression scenarios.
For synchronization enhancement (upper panels), the effectiveness of adversarial attacks shows minimal dependence on network size across all three topologies. The order parameter remains relatively constant as increases from 500 to 4000, indicating that positive perturbations maintain their synchronization-promoting capabilities regardless of system scale. This size-independence suggests that the mechanisms underlying synchronization enhancement are robust to network dimensionality.
In stark contrast, synchronization suppression (lower panels) exhibits a pronounced network size dependence across all network topologies. The effectiveness of adversarial attacks increases substantially with larger networks, as evidenced by the progressive decrease in with increasing . This trend is observed in all three network types, though BA networks show the most resistance to suppression due to their inherently higher synchronizability. The enhanced suppression effectiveness in larger networks stems from the inherent difficulty of achieving synchronization in high-dimensional systems, where the increased degrees of freedom naturally oppose collective dynamics.
Crucially, the comparison with random perturbations (crosses) confirms that this size-dependent enhancement of suppression is genuinely due to adversarial targeting rather than general noise effects. Random perturbations show little variation with network size and remain close to baseline values, demonstrating that the observed trends are specific to the strategic nature of adversarial attacks. This finding highlights a fundamental asymmetry: while larger networks become increasingly vulnerable to adversarial desynchronization, they do not become more susceptible to random disruption, underscoring the sophisticated targeting capability of adversarial perturbations.
IV.5 Real-world network applications
To demonstrate the practical applicability of our adversarial attack methodology, we applied it to two real-world networks where synchronization plays a crucial role: a power network (1138-bus system Rossi and Ahmed (2015)) and a mouse brain network Amunts et al. (2013); Rossi and Ahmed (2015). For simplicity, both networks were treated as unweighted undirected networks, and only the largest connected component was used for analysis. Figure 4 shows the order parameter as a function of coupling strength for both networks under various perturbation conditions.
Both networks exhibit behavior consistent with our findings from model networks, confirming the generalizability of adversarial attacks on network synchronization. In the power network (Fig. 4a), which shows inherently low synchronization due to its sparse structure, adversarial perturbations demonstrate particularly effective synchronization enhancement. Positive perturbations successfully elevate the network’s collective behavior across the entire coupling range, with stronger perturbations producing substantial improvements in coordination. This enhancement capability is especially valuable for power grids, where improved frequency synchronization directly translates to system stability Carrasco et al. (2006); Li and Wong (2017).
The brain network (Fig. 4b) demonstrates highly effective responses to both enhancement and suppression attacks. Positive perturbations produce strong synchronization enhancement, substantially elevating the order parameter across all coupling values. Conversely, the network shows significant vulnerability to desynchronization attacks, with negative perturbations effectively maintaining low synchronization levels even under strong coupling conditions that would normally induce robust collective behavior. This suppression capability is particularly relevant for therapeutic applications Tass (2002); Chazottes et al. (2005); Tass and Hauptmann (2007); Andrzejak et al. (2016), as pathological hypersynchronization in brain networks is associated with epileptic seizures Lehnertz et al. (2009); Jiruska et al. (2013).
These results demonstrate that adversarial attacks on network synchronization represent a genuine concern for real-world systems, while simultaneously suggesting potential applications for controlled synchronization enhancement in cases where increased coordination is desired.
V Discussion
In this study, we have demonstrated that adversarial attack strategies, originally developed for deep learning systems, can be effectively adapted to control synchronization in networked dynamical systems governed by the Kuramoto model. Our results consistently show that small, strategically applied perturbations to the phases of network nodes can dramatically enhance or suppress collective synchronization across diverse network architectures, from theoretical frameworks to practical infrastructure and biological systems.
Our results confirm that borrowing the adversarial attack concept from the deep learning domain and applying it to network synchronization control represents a significant advance. Unlike traditional synchronization control methods that typically require structural modifications Zhou and Kurths (2006); Chen et al. (2008); Zhang and Xiao (2014); Pinto and Saa (2015); Papadopoulos et al. (2017); Ghorban, Hesaam, and Sarbazi-Azad (2022); Ruangkriengsin and Porter (2023); Song et al. (2024); Menara et al. (2022) such as adding or removing links, rewiring connections, or introducing external driving forces Wang and Chen (2002); Sorrentino et al. (2007); Xiang et al. (2007); Yang et al. (2010); Lin et al. (2016); Yeung and Strogatz (1999); Rosenblum and Pikovsky (2004); Taher, Olmi, and Schöll (2019); Franci et al. (2012); Ozawa and Kori (2021); Kawamura, Nakao, and Kuramoto (2007); Esfahani, Shahbazi, and Samani (2012); Lai and Porter (2013); Tyloo, Delabays, and Jacquod (2019), our adversarial perturbation strategy achieves effective control through weak interventions. While this approach builds on well-established principles of gradient-based feedback and weak intervention strategies, our contribution lies in the specific formulation and application context that leverages the order parameter structure for distributed phase perturbations. While conventional approaches often necessitate global knowledge of network topology or significant alterations to the system architecture, our method operates by applying small amplitude perturbations () to the phases of network nodes, requiring only the computation of gradients and maintaining the original network structure intact.
It is important to distinguish our approach from related work on network perturbations, particularly Byzantine attacks on oscillator networks Tyloo (2023). While both approaches demonstrate the significant impact that strategic perturbations can have on networked oscillators, they differ fundamentally in purpose and methodology. Byzantine attacks involve malicious actors who completely control individual nodes with arbitrary input signals to disrupt or destabilize synchronization for adversarial purposes. In contrast, our gradient-based framework uses the mathematical structure of the order parameter to determine optimal perturbation directions that enhance or suppress synchronization for beneficial control objectives. Rather than arbitrary signal injection, our method employs principled optimization to achieve desired synchronization states while maintaining the integrity of the network structure. This represents a constructive application of strategic perturbation principles, transforming concepts originally associated with network vulnerabilities into tools for beneficial synchronization management.
This efficiency is particularly notable when compared to conventional control approaches. Many existing synchronization control strategies involve node selective interventions Wang and Chen (2002); Sorrentino et al. (2007); Xiang et al. (2007); Yang et al. (2010); Lin et al. (2016), structural network modifications Zhou and Kurths (2006); Chen et al. (2008); Zhang and Xiao (2014); Pinto and Saa (2015); Papadopoulos et al. (2017); Ghorban, Hesaam, and Sarbazi-Azad (2022); Ruangkriengsin and Porter (2023); Song et al. (2024), or the application of external driving forces Yeung and Strogatz (1999); Rosenblum and Pikovsky (2004); Taher, Olmi, and Schöll (2019); Franci et al. (2012); Ozawa and Kori (2021); Kawamura, Nakao, and Kuramoto (2007); Esfahani, Shahbazi, and Samani (2012); Lai and Porter (2013); Tyloo, Delabays, and Jacquod (2019) to specific system components. In contrast, our adversarial approach demonstrates that modest perturbations applied uniformly across the network can produce system wide changes in synchronization behavior. The distributed nature of this intervention offers complementary advantages for applications where uniform access to network nodes is feasible, providing an alternative pathway for synchronization control that maintains the original network architecture.
Moreover, our findings reveal the dual nature of network vulnerability and controllability: the same mechanisms that make networks susceptible to adversarial attacks also provide powerful tools for beneficial synchronization control. This insight opens new avenues for both understanding network robustness and developing efficient control strategies for complex systems where synchronized behavior is either desired or needs to be prevented.
Although we focused on the Kuramoto model in this study, the core of our attack methodology, namely the gradient computation from the order parameter, is derived purely from the mathematical definition of synchronization and is not specific to any particular oscillator model. This generality suggests that our adversarial perturbation approach can be readily extended to other oscillatory systems, such as FitzHugh–Nagumo neurons Gerster et al. (2020); Al Beattie et al. (2024), Stuart–Landau oscillators Choe et al. (2010); Selivanov et al. (2012), or any dynamical system where a meaningful order parameter can be defined. The mathematical framework we developed is thus broadly applicable to diverse classes of networked dynamical systems.
While our results demonstrate the effectiveness of adversarial perturbations for synchronization control, we acknowledge important practical considerations for real-world implementation. The current approach requires simultaneous access to all oscillator phases and the ability to perturb all nodes, which presents challenges in practical systems such as power grids or brain networks. We view our current results as establishing a theoretical foundation that opens promising avenues for future developments.
Building on this foundation, several exciting research directions emerge: developing sparse perturbation schemes that achieve effective control by targeting only strategically selected nodes, creating partial-access strategies that operate with limited system information, and designing adaptive methods that can function under realistic constraints. These developments represent natural extensions of our framework that could enable practical synchronization control in real-world networked systems while maintaining the core advantages of weak intervention and distributed control.
Additionally, the optimal balance between perturbation parameter and intervention timing requires careful consideration. In this study, values were selected empirically based on the trade-off between control effectiveness and intervention frequency, as demonstrated in our parameter sweep analysis (Fig. 2). While larger perturbations and more frequent interventions naturally enhance control effectiveness, practical applications demand minimal interference with the system. The key challenge is achieving robust synchronization control using the smallest possible perturbations applied as infrequently as possible. A deeper theoretical analysis of the attack mechanism would provide valuable insights into these parameter trade-offs and the fundamental principles underlying the observed synchronization control.
Future work should also validate our approach across different oscillator models, as well as extend the framework to directed and weighted networks Porfiri, Stilwell, and Bollt (2008); Duan, Chen, and Huang (2008); Tiberi et al. (2017), and systems with higher-order interactions Gao et al. (2023); Liu et al. (2024); Wang and Zhao (2024). These extensions would further demonstrate the broad applicability of our adversarial synchronization control paradigm.
Nonetheless, the core finding that small, strategically applied perturbations can dramatically alter collective synchronization behavior represents a significant conceptual advance. The simplicity of the approach, combined with its effectiveness across diverse network architectures, establishes a new paradigm for synchronization control that bridges deep learning and network dynamics. The broad scope for future developments, from sparse interventions to higher-order networks, underscores the fundamental importance and potential impact of this research direction.
The data and code used in this study are publicly available at https://github.com/kztakemoto/advSync.
Acknowledgements.
This study was funded by JSPS KAKENHI (grant number JP21H03545).References
- Arenas et al. (2008) A. Arenas, A. Díaz-Guilera, J. Kurths, Y. Moreno, and C. Zhou, “Synchronization in complex networks,” Physics reports 469, 93–153 (2008).
- Suykens and Osipov (2008) J. A. Suykens and G. V. Osipov, “Introduction to focus issue: synchronization in complex networks,” Chaos: An Interdisciplinary Journal of Nonlinear Science 18 (2008).
- Dörfler and Bullo (2014) F. Dörfler and F. Bullo, “Synchronization in complex networks of phase oscillators: A survey,” Automatica 50, 1539–1564 (2014).
- Acebrón et al. (2005) J. A. Acebrón, L. L. Bonilla, C. J. Pérez Vicente, F. Ritort, and R. Spigler, “The kuramoto model: A simple paradigm for synchronization phenomena,” Reviews of modern physics 77, 137–185 (2005).
- Rodrigues et al. (2016) F. A. Rodrigues, T. K. D. Peron, P. Ji, and J. Kurths, “The kuramoto model in complex networks,” Physics Reports 610, 1–98 (2016).
- Moreno and Pacheco (2004) Y. Moreno and A. F. Pacheco, “Synchronization of kuramoto oscillators in scale-free networks,” Europhysics Letters 68, 603 (2004).
- Peron et al. (2019) T. Peron, B. Messias F. de Resende, A. S. Mata, F. A. Rodrigues, and Y. Moreno, “Onset of synchronization of kuramoto oscillators in scale-free networks,” Physical Review E 100, 042302 (2019).
- Medvedev and Tang (2020) G. S. Medvedev and X. Tang, “The kuramoto model on power law graphs: synchronization and contrast states,” Journal of nonlinear science 30, 2405–2427 (2020).
- Hong, Choi, and Kim (2002) H. Hong, M.-Y. Choi, and B. J. Kim, “Synchronization on small-world networks,” Physical Review E 65, 026139 (2002).
- Grabow et al. (2010) C. Grabow, S. M. Hill, S. Grosskinsky, and M. Timme, “Do small worlds synchronize fastest?” Europhysics Letters 90, 48002 (2010).
- Nishikawa et al. (2003) T. Nishikawa, A. E. Motter, Y.-C. Lai, and F. C. Hoppensteadt, “Heterogeneity in oscillator networks: Are smaller worlds easier to synchronize?” Physical review letters 91, 014101 (2003).
- Zhou and Kurths (2006) C. Zhou and J. Kurths, “Dynamical weights and enhanced synchronization in adaptive complex networks,” Physical review letters 96, 164102 (2006).
- Chen et al. (2008) M. Chen, Y. Shang, Y. Zou, and J. Kurths, “Synchronization in the kuramoto model: A dynamical gradient network approach,” Physical Review E—Statistical, Nonlinear, and Soft Matter Physics 77, 027101 (2008).
- Zhang and Xiao (2014) Y. Zhang and R. Xiao, “Synchronization of kuramoto oscillators in small-world networks,” Physica A: Statistical Mechanics and its Applications 416, 33–40 (2014).
- Pinto and Saa (2015) R. S. Pinto and A. Saa, “Optimal synchronization of kuramoto oscillators: A dimensional reduction approach,” Physical Review E 92, 062801 (2015).
- Papadopoulos et al. (2017) L. Papadopoulos, J. Z. Kim, J. Kurths, and D. S. Bassett, “Development of structural correlations and synchronization from adaptive rewiring in networks of kuramoto oscillators,” Chaos: An Interdisciplinary Journal of Nonlinear Science 27 (2017).
- Ghorban, Hesaam, and Sarbazi-Azad (2022) S. H. Ghorban, B. Hesaam, and H. Sarbazi-Azad, “A rewiring mechanism to improve synchronization in kuramoto networks,” Journal of Statistical Mechanics: Theory and Experiment 2022, 113404 (2022).
- Ruangkriengsin and Porter (2023) T. Ruangkriengsin and M. A. Porter, “Low-dimensional behavior of a kuramoto model with inertia and hebbian learning,” Chaos: An Interdisciplinary Journal of Nonlinear Science 33 (2023).
- Song et al. (2024) Y. Song, X. Liu, D. Wang, P. Gao, and M. Xue, “Network synchronizability enhancement via adding antagonistic interactions,” Autonomous Intelligent Systems 4, 1–7 (2024).
- Acebrón and Spigler (1998) J. Acebrón and R. Spigler, “Adaptive frequency model for phase-frequency synchronization in large populations of globally coupled nonlinear oscillators,” Physical Review Letters 81, 2229 (1998).
- Taylor, Ott, and Restrepo (2010) D. Taylor, E. Ott, and J. G. Restrepo, “Spontaneous synchronization of coupled oscillator systems with frequency adaptation,” Physical Review E—Statistical, Nonlinear, and Soft Matter Physics 81, 046214 (2010).
- Gil (2021) L. Gil, “Optimally frequency-synchronized networks of nonidentical kuramoto oscillators,” Physical Review E 104, 044211 (2021).
- Ren and Zhao (2007) Q. Ren and J. Zhao, “Adaptive coupling and enhanced synchronization in coupled phase oscillators,” Physical Review E—Statistical, Nonlinear, and Soft Matter Physics 76, 016207 (2007).
- Ha, Noh, and Park (2016) S.-Y. Ha, S. E. Noh, and J. Park, “Synchronization of kuramoto oscillators with adaptive couplings,” SIAM Journal on Applied Dynamical Systems 15, 162–194 (2016).
- Cho, Dong, and Ha (2023) H. Cho, J.-G. Dong, and S.-Y. Ha, “Interplay of inertia and adaptive couplings in the emergent dynamics of kuramoto ensemble,” Journal of Differential Equations 360, 523–571 (2023).
- Wang and Li (2011) H. Wang and X. Li, “Synchronization and chimera states of frequency-weighted kuramoto-oscillator networks,” Physical Review E—Statistical, Nonlinear, and Soft Matter Physics 83, 066214 (2011).
- Rossi et al. (2023) K. L. Rossi, R. C. Budzinski, B. R. Boaretto, L. E. Muller, and U. Feudel, “Shifts in global network dynamics due to small changes at single nodes,” Physical Review Research 5, 013220 (2023).
- Menara et al. (2022) T. Menara, G. Baggio, D. Bassett, and F. Pasqualetti, “Functional control of oscillator networks,” Nature communications 13, 4721 (2022).
- Wang and Chen (2002) X. F. Wang and G. Chen, “Pinning control of scale-free dynamical networks,” Physica A: Statistical Mechanics and its Applications 310, 521–531 (2002).
- Sorrentino et al. (2007) F. Sorrentino, M. Di Bernardo, F. Garofalo, and G. Chen, “Controllability of complex networks via pinning,” Physical Review E—Statistical, Nonlinear, and Soft Matter Physics 75, 046103 (2007).
- Xiang et al. (2007) L. Xiang, Z. Liu, Z. Chen, F. Chen, and Z. Yuan, “Pinning control of complex dynamical networks with general topology,” Physica A: Statistical Mechanics and its Applications 379, 298–306 (2007).
- Yang et al. (2010) L. Yang, X. Wang, Y. Li, and Z. Sheng, “On the pinning strategy of complex networks,” Europhysics Letters 92, 48002 (2010).
- Lin et al. (2016) W. Lin, H. Fan, Y. Wang, H. Ying, and X. Wang, “Controlling synchronous patterns in complex networks,” Physical Review E 93, 042209 (2016).
- Yeung and Strogatz (1999) M. S. Yeung and S. H. Strogatz, “Time delay in the kuramoto model of coupled oscillators,” Physical review letters 82, 648 (1999).
- Rosenblum and Pikovsky (2004) M. Rosenblum and A. Pikovsky, “Delayed feedback control of collective synchrony: An approach to suppression of pathological brain rhythms,” Physical Review E—Statistical, Nonlinear, and Soft Matter Physics 70, 041904 (2004).
- Taher, Olmi, and Schöll (2019) H. Taher, S. Olmi, and E. Schöll, “Enhancing power grid synchronization and stability through time-delayed feedback control,” Physical Review E 100, 062306 (2019).
- Franci et al. (2012) A. Franci, A. Chaillet, E. Panteley, and F. Lamnabhi-Lagarrigue, “Desynchronization and inhibition of kuramoto oscillators by scalar mean-field feedback,” Mathematics of Control, Signals, and Systems 24, 169–217 (2012).
- Ozawa and Kori (2021) A. Ozawa and H. Kori, “Feedback-induced desynchronization and oscillation quenching in a population of globally coupled oscillators,” Physical Review E 103, 062217 (2021).
- Kawamura, Nakao, and Kuramoto (2007) Y. Kawamura, H. Nakao, and Y. Kuramoto, “Noise-induced turbulence in nonlocally coupled oscillators,” Physical Review E—Statistical, Nonlinear, and Soft Matter Physics 75, 036209 (2007).
- Esfahani, Shahbazi, and Samani (2012) R. K. Esfahani, F. Shahbazi, and K. A. Samani, “Noise-induced synchronization in small world networks of phase oscillators,” Physical Review E—Statistical, Nonlinear, and Soft Matter Physics 86, 036204 (2012).
- Lai and Porter (2013) Y. M. Lai and M. A. Porter, “Noise-induced synchronization, desynchronization, and clustering in globally coupled nonidentical oscillators,” Physical Review E—Statistical, Nonlinear, and Soft Matter Physics 88, 012905 (2013).
- Tyloo, Delabays, and Jacquod (2019) M. Tyloo, R. Delabays, and P. Jacquod, “Noise-induced desynchronization and stochastic escape from equilibrium in complex networks,” Physical Review E 99, 062213 (2019).
- Tyloo (2023) M. Tyloo, “Assessing the impact of byzantine attacks on coupled phase oscillators,” Journal of Physics: Complexity 4, 045005 (2023).
- Yuan et al. (2019) X. Yuan, P. He, Q. Zhu, and X. Li, “Adversarial examples: Attacks and defenses for deep learning,” IEEE transactions on neural networks and learning systems 30, 2805–2824 (2019).
- Hirano and Takemoto (2020) H. Hirano and K. Takemoto, “Simple iterative method for generating targeted universal adversarial perturbations,” Algorithms 13, 268 (2020).
- Hirano, Koga, and Takemoto (2020) H. Hirano, K. Koga, and K. Takemoto, “Vulnerability of deep neural networks for detecting covid-19 cases from chest x-ray images to universal adversarial attacks,” Plos one 15, e0243963 (2020).
- Hirano, Minagi, and Takemoto (2021) H. Hirano, A. Minagi, and K. Takemoto, “Universal adversarial attacks on deep neural networks for medical image classification,” BMC medical imaging 21, 1–13 (2021).
- Chakraborty et al. (2021) A. Chakraborty, M. Alam, V. Dey, A. Chattopadhyay, and D. Mukhopadhyay, “A survey on adversarial attacks and defences,” CAAI Transactions on Intelligence Technology 6, 25–45 (2021).
- Akhtar et al. (2021) N. Akhtar, A. Mian, N. Kardan, and M. Shah, “Advances in adversarial attacks and defenses in computer vision: A survey,” IEEE Access 9, 155161–155196 (2021).
- Koga and Takemoto (2022) K. Koga and K. Takemoto, “Simple black-box universal adversarial attacks on deep neural networks for medical image classification,” Algorithms 15, 144 (2022).
- Chiyomaru and Takemoto (2022) K. Chiyomaru and K. Takemoto, “Adversarial attacks on voter model dynamics in complex networks,” Physical Review E 106, 014301 (2022).
- Chiyomaru and Takemoto (2023) K. Chiyomaru and K. Takemoto, “Mitigation of adversarial attacks on voter model dynamics by network heterogeneity,” Journal of Physics: Complexity 4, 025009 (2023).
- Mizutaka (2023) S. Mizutaka, “Crossover phenomenon in adversarial attacks on voter model,” Journal of Physics: Complexity 4, 035009 (2023).
- Takemoto (2024) K. Takemoto, “Steering cooperation: Adversarial attacks on prisoner’s dilemma in complex networks,” Physica A: Statistical Mechanics and its Applications 655, 130214 (2024).
- Ninomiya et al. (2025) M. Ninomiya, G. Ichinose, K. Chiyomaru, and K. Takemoto, “Mitigating opinion polarization in social networks using adversarial attacks,” Scientific Reports 15, 9033 (2025).
- Goodfellow, Shlens, and Szegedy (2015) I. J. Goodfellow, J. Shlens, and C. Szegedy, “Explaining and harnessing adversarial examples,” in 3rd International Conference on Learning Representations, ICLR 2015, San Diego, CA, USA, May 7-9, 2015, Conference Track Proceedings, edited by Y. Bengio and Y. LeCun (2015).
- Albert and Barabási (2002) R. Albert and A.-L. Barabási, “Statistical mechanics of complex networks,” Reviews of modern physics 74, 47 (2002).
- Takemoto and Oosawa (2012) K. Takemoto and C. Oosawa, “Introduction to complex networks: Measures, statistical properties, and models,” in Statistical and Machine Learning Approaches for Network Analysis (John Wiley & Sons, Ltd, 2012) Chap. 2, pp. 45–75.
- Barabási and Albert (1999) A.-L. Barabási and R. Albert, “Emergence of scaling in random networks,” Science 286, 509–512 (1999).
- Watts and Strogatz (1998) D. J. Watts and S. H. Strogatz, “Collective dynamics of ‘small-world’networks,” Nature 393, 440–442 (1998).
- Hirano and Takemoto (2019) H. Hirano and K. Takemoto, “Difficulty in inferring microbial community structure based on co-occurrence network approaches,” BMC bioinformatics 20, 1–14 (2019).
- Rossi and Ahmed (2015) R. A. Rossi and N. K. Ahmed, “The network data repository with interactive graph analytics and visualization,” in AAAI (2015).
- Amunts et al. (2013) K. Amunts, C. Lepage, L. Borgeat, H. Mohlberg, T. Dickscheid, M.-É. Rousseau, S. Bludau, P.-L. Bazin, L. B. Lewis, A.-M. Oros-Peusquens, et al., “Bigbrain: an ultrahigh-resolution 3d human brain model,” science 340, 1472–1475 (2013).
- Carrasco et al. (2006) J. M. Carrasco, L. G. Franquelo, J. T. Bialasiewicz, E. Galván, R. C. PortilloGuisado, M. M. Prats, J. I. León, and N. Moreno-Alfonso, “Power-electronic systems for the grid integration of renewable energy sources: A survey,” IEEE Transactions on industrial electronics 53, 1002–1016 (2006).
- Li and Wong (2017) B. Li and K. M. Wong, “Optimizing synchronization stability of the kuramoto model in complex networks and power grids,” Physical Review E 95, 012207 (2017).
- Tass (2002) P. A. Tass, “Desynchronization of brain rhythms with soft phase-resetting techniques,” Biological cybernetics 87, 102–115 (2002).
- Chazottes et al. (2005) J.-R. Chazottes, B. Fernandez, Y. Maistrenko, O. Popovych, and P. Tass, “Desynchronization and chaos in the kuramoto model,” Dynamics of coupled map lattices and of related spatially extended systems , 285–306 (2005).
- Tass and Hauptmann (2007) P. A. Tass and C. Hauptmann, “Therapeutic modulation of synaptic connectivity with desynchronizing brain stimulation,” International Journal of Psychophysiology 64, 53–61 (2007).
- Andrzejak et al. (2016) R. G. Andrzejak, C. Rummel, F. Mormann, and K. Schindler, “All together now: Analogies between chimera state collapses and epileptic seizures,” Scientific reports 6, 23000 (2016).
- Lehnertz et al. (2009) K. Lehnertz, S. Bialonski, M.-T. Horstmann, D. Krug, A. Rothkegel, M. Staniek, and T. Wagner, “Synchronization phenomena in human epileptic brain networks,” Journal of neuroscience methods 183, 42–48 (2009).
- Jiruska et al. (2013) P. Jiruska, M. De Curtis, J. G. Jefferys, C. A. Schevon, S. J. Schiff, and K. Schindler, “Synchronization and desynchronization in epilepsy: controversies and hypotheses,” The Journal of physiology 591, 787–797 (2013).
- Gerster et al. (2020) M. Gerster, R. Berner, J. Sawicki, A. Zakharova, A. Škoch, J. Hlinka, K. Lehnertz, and E. Schöll, “Fitzhugh–nagumo oscillators on complex networks mimic epileptic-seizure-related synchronization phenomena,” Chaos: An Interdisciplinary Journal of Nonlinear Science 30 (2020).
- Al Beattie et al. (2024) B. Al Beattie, P. Feketa, K. Ochs, and H. Kohlstedt, “Criticality in fitzhugh-nagumo oscillator ensembles: Design, robustness, and spatial invariance,” Communications Physics 7, 46 (2024).
- Choe et al. (2010) C.-U. Choe, T. Dahms, P. Hövel, and E. Schöll, “Controlling synchrony by delay coupling in networks: from in-phase to splay and cluster states,” Physical Review E—Statistical, Nonlinear, and Soft Matter Physics 81, 025205 (2010).
- Selivanov et al. (2012) A. A. Selivanov, J. Lehnert, T. Dahms, P. Hövel, A. L. Fradkov, and E. Schöll, “Adaptive synchronization in delay-coupled networks of stuart-landau oscillators,” Physical Review E—Statistical, Nonlinear, and Soft Matter Physics 85, 016201 (2012).
- Porfiri, Stilwell, and Bollt (2008) M. Porfiri, D. J. Stilwell, and E. M. Bollt, “Synchronization in random weighted directed networks,” IEEE Transactions on Circuits and Systems I: Regular Papers 55, 3170–3177 (2008).
- Duan, Chen, and Huang (2008) Z. Duan, G. Chen, and L. Huang, “Synchronization of weighted networks and complex synchronized regions,” Physics Letters A 372, 3741–3751 (2008).
- Tiberi et al. (2017) L. Tiberi, C. Favaretto, M. Innocenti, D. S. Bassett, and F. Pasqualetti, “Synchronization patterns in networks of kuramoto oscillators: A geometric approach for analysis and control,” in 2017 IEEE 56th Annual Conference on Decision and Control (CDC) (IEEE, 2017) pp. 481–486.
- Gao et al. (2023) Z. Gao, D. Ghosh, H. Harrington, J. Restrepo, and D. Taylor, “Dynamics on networks with higher-order interactions,” Chaos: An Interdisciplinary Journal of Nonlinear Science 33 (2023).
- Liu et al. (2024) H. Liu, J. Zhou, B. Li, M. Huang, J.-a. Lu, and D. Shi, “Synchronization on higher-order networks,” Europhysics Letters 145, 61001 (2024).
- Wang and Zhao (2024) Y. Wang and Y. Zhao, “Synchronization of directed higher-order networks via pinning control,” Chaos, Solitons & Fractals 185, 115062 (2024).